OpenStudy (anonymous):

Help please will give medal Find the simple interest. Then find the total of principal plus interest Principal=$200 Interest rate=4% Time=2 years 2. Principal=$870 Interest rate=3% Time=9 months

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@ayee2000

OpenStudy (campbell_st):

you can use the formula \[I = \frac{Prt}{100}\] P = Principal, r = interest rate and t = time periods. substitute and calculate the answer. then add the principal and interest to get the total value. in question 2, I'd assume the interest is per annum or per year. Which means the time of 9 months needs to be changes to years... 9 months = 0.75 years hope it helps

OpenStudy (anonymous):

Thanks @campbell_st

OpenStudy (anonymous):

@campbell_st i tried to calculate it i cant i tried i asked my mom she didn't know either so can i get some more help then i will see if i can get u some more medals

OpenStudy (anonymous):

@DanJS i need some help

OpenStudy (campbell_st):

question 1 looks like \[I = \frac{200 \times 4 \times 2}{100}\] then total = 200 + the value for I you calculated above